OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: x + y = 3 x = y – 7 A. The solution is (–2, 5). B. The solution is (4, 1). C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (phi):

x = y – 7 means x "is the same" as y-7 replace x in the first equation with y-7. you get a new equation. can you write down the new equation?

OpenStudy (anonymous):

Substitute 'y - 7' for 'x' in the 1st equation: x + y = 3 y - 7 + y = 3 Combine like terms: 2y - 7 = 3 Add 7 to both sides, what's 3 + 7?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

@iGreen. is the answer A?